OpenStudy (anonymous):

The sum of the digits of a three digit number is 19. The second digit is one more than the first digit. The third digit is two more than the second digit. Find the number. 568 345 124 890 @Joel_the_boss

hartnn (hartnn):

which of those 4 choices give you sum of digits as 19 ?

hartnn (hartnn):

5+6+8 = .. ? 3+4+5 = .. ? 1+2+4 = ...? 8+9+0 =...? the one which gives you 19 is your answer :)

OpenStudy (anonymous):

Hahaha this is why math exams are pretty easy when there are choices. You could just try them all and choose whichever satisfy the give.n